A ray of light is incident at an angle 52° on the face of equilateral prism and gets through 34°.find the angle of emergence

Grade:11

2 Answers

Saurabh Koranglekar
3 years ago
For an equilateral prism is -

A = 60°

Angle of emergence of the light ray is calculated by formula -

i + e = A + d

52 + e = 60 + 34

Therefore, the angle of emergence of the ray is 42°.

Tahura Fatima sheikh
2 years ago
A ray light is incident at an angle 52 on the face of equivalent prism and gets through 34,find angle of emergence